# LoRA Users Beware: A Few Spurious Tokens Can Manipulate Your Finetuned Model

### v1

@@ 3.2 Spurious Token Injection
Figure 3: Code demonstrating a basic use of our library to inject a randomly generated date token into a basic sentence. For further examples using the code library refer to fig. 7.
A similar HTMLInjection modifier allows for injecting HTML tags (details can be found in section A.2). A code example of SSTI using the library can be found at fig. 3, with further examples at section A.3.4. Examples of what these various SSTI strategies look like can be found in fig. 4 (see section A.3 for further examples, with sections A.3.1, A.3.2 and A.3.3 showing the different token types).
Figure 4: Examples of spurious token injection (SSTI) strategies. Injected tokens are highlighted in red. Top: Original sentence without corruption. Next rows: A single token (date) is inserted at the beginning; multiple random tokens are injected at random positions; and HTML tags are inserted at the end. These patterns mimic real-world artifacts and are sufficient to steer model predictions. Our full evaluation systematically varies token type, number, and injection location (start, end, random). Additional examples in section A.3.
3.3 Procedure
We used LoRA to fine-tune a range of models across diverse datasets to evaluate the effect of spurious token injection (SSTI) on model robustness. Our experiments included five models from three major families—Snowflake Arctic [13] (arctic-embed-xs (22M), arctic-embed-l (335M)), Apple OpenELM [22] (openelm-270m (270M), openelm-3b (3B)), and Meta-LLaMA-3 [2] (llama-3-8b (8B))—covering a range of model sizes. To assess generalization, we evaluated on four datasets: IMDB [19], Financial Classification [23], CommonSenseQA [31], and Bias in Bios [9].
Each model was fine-tuned using LoRA (Hugging Face’s PEFT implementation [20]) with ranks of 1, 16, 32, and 64, on frozen pretrained weights. Training hyperparameters were scaled to model size: smaller models (under 1B parameters) used a per-device batch size of 16, weight decay of $1e^{-5}$, and a learning rate of $1e^{-4}$, while larger models used a per-device batch size ranging from 1 to 14 to accommodate memory constraints and dataset size. They were all trained until convergence, so the number of training steps differed. For full software and hardware details, including GPU type and infrastructure, see section A.1.
For SSTI, we used a controlled spurious token injection framework. All injections were added only to samples with a particular class label. We systematically varied the following. Proportion of samples injected: 0%, 25%, 50%, 75%, 100%. Token proportion: 1 token, 5% of each injected sample’s original tokens, or 10%. Token type: dates, countries, or HTML tags. SSTI location: beginning, end, or random. Each configuration was evaluated on both a clean test set and a matched spurious test set, using the same token injection parameters applied during training. This dual-evaluation framework allows us to assess both real-world deployment behavior (with latent spurious correlations) and clean generalization performance. For an overview of the injection procedure and examples of injected tokens, see section 3.2 and section A.3.
4 LoRA Feeds on Spurious Tokens
This section explores how and when LoRA-finetuned models become vulnerable to spurious token injection (SSTI). In section 4.1, we show that even minimal corruption—just a single token per prompt— is sufficient to control model predictions. Section 4.2 demonstrates the suprising finding that increasing LoRA rank under Light SSTI amplifies this vulnerability. Section 4.3 reveals a reversal: under Aggressive SSTI, higher ranks help recover robustness by attending to non-spurious features. Together, these results expose a non-monotonic relationship between LoRA capacity and robustness.
In section 4.4, we show that SSTI is able to control the model’s behaviour regardless of where the spurious token is injected or what form it takes. In section 4.5 we show that using a larger model or finetuning for longer does not solve this. Finally, in section 4.6, we go under the hood to reveal how this shortcut reliance surfaces in model internals, showing that attention entropy provides a useful diagnostic for detecting spurious behavior.
4.1 A Single Token Can Manipulate the Model
We begin our analysis with the Light SSTI setting, where only a single spurious token is injected per prompt and correlated with a specific class. We ask the question: Is such minimal corruption sufficient to alter model behavior?
As shown in LoRA Users Beware: A Few Spurious Tokens Can Manipulate Your Finetuned Model, the answer is yes. When training samples are injected with a single token associated with a target class, the model trained under this corruption overwhelmingly predicts that class at test time—regardless of input content. For example, injecting a class 0-associated token results in the model assigning nearly all test samples to class 0. In contrast, the base model distributes predictions more evenly across classes. This result demonstrates that even minimal, single-token corruption is sufficient to deterministically control model outputs.
Figure 5 (left) shows a surprising trend: under Light SSTI, increasing LoRA rank leads to a widening gap between performance on clean and spurious test sets. Clean accuracy remains mostly flat, while spurious-set performance improves sharply—indicating that the model has learned to rely on the injected token rather than generalizing from meaningful task features. This pattern becomes more evident in fig. 5 (right), which plots the difference in accuracy between spurious and clean evaluations across ranks and injection proportions. Even when only 25–50% of training samples contain the spurious token, the performance gap grows with rank. The effect is particularly pronounced at 50% and above, suggesting that under light SSTI, higher-rank adapters are more prone to overfitting to spurious correlations (higher LoRA capacity increases the model’s tendency to exploit shortcut correlations, even when those correlations are sparse).
These results extend the finding from section 4.1: not only is minimal corruption sufficient to steer predictions, but this vulnerability is amplified as LoRA rank increases. In section 4.3, we examine whether this trend persists under more aggressive forms of SSTI—where spurious signals are more dominant and more frequent.
4.3 Aggressive SSTI: Greater Rank = Greater Robustness
In section 4.2, we showed that under Light SSTI, increasing LoRA rank exacerbates a model’s reliance on spurious signals. But what happens when the corruption is no longer minimal?
To explore this, we performed the same experiments under a more aggressive SSTI setting—where 50% of training samples are injected with spurious tokens amounting to 10% of each sample’s token count. Surprisingly, under this regime, we observe a reversal of the earlier trend: higher LoRA ranks now begin to improve robustness, rather than hurt it. Figure 6 (left) illustrates this shift. Unlike the Light SSTI case, the gap between clean and spurious evaluation accuracy narrows as LoRA rank increases. This suggests that higher-capacity adapters are better equipped to reconcile conflicting training signals, and recover generalization in the face of strong spurious signals. Figure 6 (right) provides a more granular view, showing balanced accuracy across LoRA ranks on clean vs. spurious test sets. While low-rank models continue to overfit the spurious tokens, higher-rank models achieve more balanced performance—no longer relying entirely on shortcut features, but instead recovering aspects of the true task signal.
Together, these results highlight a key insight: the relationship between LoRA capacity and robustness is non-monotonic. When spurious signals are weak, low-rank adapters act as a regularizer by limiting memorization. But as spurious signals become more dominant, higher ranks enable the model to better interpolate between noisy and clean supervision—improving test-time alignment. We observed similar trends across other datasets and model scales, as detailed in section A.6. In the next section (section 4.4), we analyze whether this behavior of SSTI controlling model behavior depends on token location and type, confirming that these trends generalizes across artifact structures.
4.4 Token Location and Type Don’t Matter
Building on the patterns established in section 4.2 and section 4.3, we now ask whether LoRA’s susceptibility to spurious tokens depends on the form or position of those tokens—i.e., whether the vulnerability is tied to specific injection artifacts or represents a more general failure mode. To probe this, we conducted two sets of controlled experiments. First, we varied the position of the injected token—beginning, end, or random—while keeping all other factors constant. Second, we varied the type of injected token (e.g., dates, country names, HTML tags).
Although minor variations exist within our trends, the overarching behavior remains consistent (as seen in table 8), suggesting that the observed behavior is not tied to any specific artifact structure or token position. Rather, it reflects a broader vulnerability of LoRA-based models to systematic dataset perturbations. Additional experiments on varying token injection locations and types are provided in sections A.8 and A.7. Further results involving variations of the "diversity" of tokens. Results of these are provided in section A.9. Together, these findings show that the shortcut reliance observed in the previous sections is not brittle—it persists across variations in token form and position. In section 4.5 we investigate whether this behavior persists when using a larger model or finetuning for longer.
Table 2: Accuracy degradation (↓, in percentage points) across two perturbation dimensions—injection location and token type—for snowflake-arctic-embed-l on the IMDB dataset. Results are shown for both Light and Aggressive SSTI (with 50% samples injected). An outlier for the light SSTI trend with date tokens, but is consistent across locations. Becomes consistent with the light SSTI trend: higher rank amplifies susceptibility for other token types, for date and HTML tokens. Fully consistent for aggressive SSTI: high rank improves robustness. For all cases, SSTI controls the behavior of the model.
SSTI Rank Injection Location Token Type
Beg. End Rand Date Country HTML
Light 1 4.14 4.21 4.24 4.21 0.67 0.74
16 4.14 4.07 4.09 4.07 2.07 1.79
32 4.02 3.82 3.91 3.82 2.91 2.45
64 3.80 3.62 3.59 3.62 3.00 2.84
Agg. 1 11.64 11.54 11.66 11.54 8.25 9.91
16 4.62 4.58 4.58 4.58 4.40 4.72
32 4.35 4.25 4.36 4.25 4.16 4.54
64 4.09 3.95 4.03 3.95 3.92 4.26
4.5 Larger Models and Longer Finetuning Does Not Help
In section 4.4 we showed that SSTI can control model behaviour regardless of the location and type of the injected tokens. In this section, we assess whether using a larger model or finetuning for longer can help. To do this, we conducted two additional experiments. One with mistralai/Mistral-Small-24B-Base-2501 [1], a 24B parameter model with extensive pretraining. The other using snowflake-arctic-embed-xs, varying the number of training steps (500, 5000, 30000).
Due to hardware constraints, we were only able to run the large model experiment for 7,500 training steps. Nonetheless, the results were striking: even this larger-parameter model exhibited substantial degradation under SSTI. This can be seen in table 3. The ablation on the number of training steps paints an equally striking picture. Training for longer does not appear to remove the effects of SSTI (see table 4). Further, table 4 also shows that the behavior from section 4.3, with a higher LoRA rank increasing robustness under aggressive SSTI, continues regardless of the number of training steps.
In section 4.6, we investigate how this vulnerability manifests internally, and whether it can be detected from the model’s attention patterns.
Model Parameters Accuracy Degradation (@ 7,500 steps)
mistralai/Mistral-Small-24B-Base-2501 24B 12.256 (pp)
Table 4: Difference in balanced accuracy between spurious and clean evaluation sets (accuracy degradation in pp) across LoRA ranks for agressive SSTI on snowflake-arctic-embed-xs and IMDB. Fine-tuning for different amounts of steps. SSTI controls model behavior despite longer training.
Number of Training Steps Rank
1 16 32 64
500 20.14 8.26 7.71 6.97
5,000 6.95 5.07 4.72 4.26
30,000 5.27 4.46 4.50 4.34
4.6 Under the hood: Detect Spurious Tokens via Attention Entropy
In the previous sections (section 4.1, section 4.2, section 4.3, section 4.4, and section 4.5), we saw when and how LoRA-finetuned models become susceptible to spurious token injection (SSTI), showing that even minimal corruption can steer predictions, and that LoRA rank modulates this vulnerability. In this section, we turn inward to examine how this vulnerability manifests inside the model. Specifically, we ask: does shortcut reliance leave a detectable trace in the model’s attention patterns?
To probe this, we visualize token-level attention distributions using the TAHV library [37], focusing on the smallest model in our suite, snowflake-arctic-embed-xs, on the IMDB dataset (table 5). We compare samples with and without injected spurious tokens and observe that when SSTI is present, attention becomes sharply concentrated on specific tokens. To quantify this concentration, we compute the Shannon entropy over token-level attention scores. Intuitively, a model relying on a spurious shortcut should exhibit lower entropy, as its attention collapses onto the injected token. Indeed, across a variety of settings, we consistently observe lower entropy in the spurious class compared to the non-spurious class. While the absolute difference varies, a reliable pattern emerges: in all cases, the entropy for spurious samples remains below 95% of that of the non-spurious samples.
This suggests a practical heuristic for diagnosing SSTI: if the attention entropy for one class is consistently below 95% of the other, the dataset may exhibit spurious correlations, and should be investigated further. Importantly, this visualization-based diagnostic makes model vulnerability both observable and quantifiable. For extended results across LoRA ranks and injection intensities, see section A.10 and its corresponding tables tables 9, 10, 11 and 12.
Table 5: Token-level attention visualizations for samples with (top) and without (bottom) SSTI, using LoRA rank 1, 10% token injection, and 50% spurious sample rate on snowflake-arctic-embed-xs (Head 0). Cropped samples to conserve space, for full visualization of samples refer to table 10. When SSTI is present, attention is more concentrated, resulting in lower entropy ($\approx 6.90$ vs. $\approx 7.60$). SSTI doesn’t just influence predictions—it warps what the model pays attention to.
Tokens Attended To Category Entropy
5 Conclusion
In this work, we expose a critical vulnerability in parameter-efficient finetuning (PEFT), demonstrating that even minimal spurious token injection can drastically influence model behavior. Through a systematic evaluation across multiple LoRA configurations, model families, and datasets, we conclude the following. Single-token injection suffices: Injecting just a single token per prompt is sufficient to steer model predictions. LoRA rank amplifies or mitigates vulnerability depending on context: Under light SSTI, higher LoRA ranks increase reliance on spurious tokens; under aggressive SSTI, they help restore robustness. Location and Type of injected token don’t matter: SSTI is able to control model behavior regardless of where tokens are injected (beginning, end, or random) and what kind of token is used (dates, HTML tags, or country names). Larger Models and Longer Finetuning Does Not Help: Using a larger model or finetuning for longer does not stop SSTI from controlling model behavior. Attention entropy can help detect SSTI reliance: Spurious tokens sharply reduce entropy in attention heatmaps, indicating model over-reliance on a narrow set of features. We find that if the entropy of spurious samples drops consistently below 95% of that of non-spurious samples, it may indicate SSTI is present.
Taken together, our results expose a fundamental tradeoff between the efficiency of modern PEFT methods and their robustness to subtle dataset corruptions. We urge practitioners to look beyond clean benchmark performance and treat robustness evaluation as a core component of the finetuning pipeline—particularly when using crowd-sourced, templated, or externally provided data, where spurious artifacts may be common and hard to detect.
Future directions.
While our experiments focus on classification-style tasks, an open question remains: how do similar spurious signals manifest in generative settings like next-token prediction? Unlike classification, the boundary between signal and shortcut is blurrier in generation, making it harder to define and detect spurious patterns. We believe this represents a rich direction for future work, and encourage the community to build analogous SSTI-style tests for such language modeling. Finally, while our SSTI method operates on the dataset side, it is equally important to investigate vulnerabilities from the model side. A malicious actor could release a pretrained model embedded with hidden control triggers—tokens that appear benign but drastically influence downstream behavior post-finetuning.
We also release an SSTI injection toolkit to help researchers test their own pipelines, as well as facilitate future research into different types of corruptions (https://github.com/pradyut3501/spurious_corr).
Appendix A Technical Appendices and Supplementary Material
Here we share some supplementary figures and information related to our experiments and results. These results support our findings across multiple models, datasets, and also provide further findings. It is divided into roughly eleven sections in which the following can be found: Information on the model and resources used (section A.1), further examples covering everything to do with SSTI, its code, and injection (sections A.2 and A.3) , an overview of the conditional token entropy on clean datasets (section A.4), a continuation of our exploration of higher LoRA ranks being more robust under agressive SSTI (section A.5), an expansion of our observations across locations (section A.7), a similar expansion but across SSTI token types (section A.8), some additional examples of trying to recognize SSTI (section A.10), and finally what the loss looks like when training our models (section A.11).
A.1 Resources Used
Table 6: Information on Datasets Used
Name Number of Categories Train/Test Size (in thousands)
IMDB [19] 2 25 / 25
Financial Classification [23] 3 4.55 / 0.506
Bias in Bios [9] 28 257 / 99.1
Common Sense [31] 5 9.74 / 1.22
Table 7: Information on Models Used
[…]
snowflake-arctic-embed-xs [13] 22M $12min$ / $3m$ / $2hm$ / $5m$
[…]
OpenELM-270M [22] 270M $2hrs$ / $13m$ /  $20h20m$ / $48m$
OpenELM-3B [22] 3B $1d2hrs$ / $3hrs$ / N/A / $1h5m$
Meta-Llama-3-8B [2] 8B $11hrs$ / $51m$ / N/A / $3h12m$
Each model was fine-tuned using LoRA with ranks of 1, 16, 32, and 64, on frozen pretrained weights. Training hyperparameters were scaled to model size: smaller models (under 1B parameters) used a per-device batch size of 16, 500 training steps, weight decay of $1e^{-5}$, and a learning rate of $1e^{-4}$, while larger models used a per-device batch size ranging from 2 to 14 to accommodate memory constraints and dataset sizes. These different batch sizes sometimes changed the amount of time steps the model was trained for but we took this as a good opportunity, allowing us to test different time steps as well.
A.2 HTML Injection Explanation
A.3 Spurious Token Injection Examples
A.3.1 Dates SSTI Examples
Injection Strategy Injected Example (Dates)
A.3.2 HTML SSTI Examples
SSTI Example Table 2: SSTI injection examples using HTML tokens (sample prompt from Financial Classification dataset). Injection varies by position and token count (for details on HTML Injection methodology, see section A.2). Injected tokens are shown in red.
Injection Strategy Injected Example (HTML)
A.3.3 Countries SSTI Examples
Injection Strategy Injected Example (Countries)
A.3.4 SSTI Code Examples
One of the central contributions of this paper is the release of a plug-and-play framework for injecting spurious corruptions into Hugging Face datasets. This toolkit is designed to make it easy for practitioners and researchers to test model robustness under spurious correlations and to facilitate future work on additional corruption strategies. The codebase is available at https://github.com/pradyut3501/spurious_corr.
Section 3.2 details the core components of the framework, including the Modifier base class, the ItemInjection and HTMLInjection implementations, and the spurious_transform function. The latter enables the creation of controlled spurious correlations by selectively applying a given modifier to a user-specified proportion of training samples associated with a target label. In this section, we walk through a few basic code examples that demonstrate the core functionality of the framework. Further examples can be found at https://github.com/pradyut3501/spurious_corr/tree/main/examples/
@@ A.4 Entropy
A.5 LoRA Continues to Feed on SSTI
We provide additional figures illustrating how LoRA-based finetuning allows spurious token injection (SSTI) to control and hijack a model, resulting in accuracy degradation. Specifically, we ablate over the proportion of injected tokens—starting with a single token and scaling up to 10% of input tokens. These examples confirm that the vulnerability persists and intensifies as the amount of SSTI increases. Additional figures throughout the appendix reinforce this finding and highlight how SSTI continues to dominate model behavior across settings: see section A.6, section A.7, section A.8, and section A.9.
@@ A.5 LoRA Continues to Feed on SSTI
A.6 Aggresive SSTI: Robustness Through LoRA Rank
We further analyze the trend reversal across different models and datasets. Under Aggressive SSTI, we consistently observe that higher LoRA ranks begin to outperform lower ones—reversing the earlier pattern seen under Light SSTI, where smaller ranks showed less susceptibility. This reversal suggests that when spurious signals are strong and frequent, larger adapters help models recover robustness by attending to more meaningful patterns.
While some exceptions exist, they typically occur when the model underperforms on the downstream task. In such cases, extending training time often closes the performance gap or restores the robustness trend in favor of higher LoRA ranks.
@@ A.6 Aggresive SSTI: Robustness Through LoRA Rank
OpenELM-3B NA NA NA NA
Meta-Llama-3-8B NA NA NA NA
A.7 Trends Across Locations
@@
(b) Difference in balanced accuracy (↓) between spurious and clean test sets across LoRA ranks, under aggressive SSTI. Each curve corresponds to a different LoRA rank. LoRA rank amplifies resistance to spurious correlations when injection occurs at a random location in the samples.
(c) Difference in balanced accuracy (↓) between spurious and clean test sets across LoRA ranks, under aggressive SSTI. Each curve corresponds to a different LoRA rank. LoRA rank amplifies resistance to spurious correlations when injection occurs at a random location in the samples.
@@
A.8 Other Spurious Token Types
@@
A.9 Impact of token diversity
In this section, we look at how token diversity impacts our results. We ablate on date tokens due to the higher ceiling of unique tokens available, allowing us to test a wider range of token diversity that would be possible with tokens or countries.
@@ A.9 Impact of token diversity
A.10 Further Examples for Recognizing SSTI

### v2

@@ A.4.6 SSTI Code Examples
Figure 5: Code demonstrating a basic use of our library to inject a randomly generated date token into a basic sentence. For further examples using the code library refer to the rest of examples in this fig.˜6
@@ A.5 Entropy
A.6 LoRA Continues to Feed on SSTI
We provide additional figures illustrating how LoRA-based finetuning allows spurious token injection (SSTI) to control and hijack a model, resulting in accuracy degradation. Specifically, we ablate over the proportion of injected tokens—starting with a single token and scaling up to 10% of input tokens. These examples confirm that the vulnerability persists and intensifies as the amount of SSTI increases. Additional figures throughout the appendix reinforce this finding and highlight how SSTI continues to dominate model behavior across settings: see section˜A.7, section˜A.8, section˜A.9, and section˜A.10.
@@ A.6 LoRA Continues to Feed on SSTI
A.7 Aggresive SSTI: Robustness Through LoRA Rank
We further analyze the trend reversal across different models and datasets. Under Aggressive SSTI, we consistently observe that higher LoRA ranks begin to outperform lower ones—reversing the earlier pattern seen under Light SSTI, where smaller ranks showed less susceptibility. This reversal suggests that when spurious signals are strong and frequent, larger adapters help models recover robustness by attending to more meaningful patterns. This can be seen in fig.˜9 (left). A more granular view showing these trends can be found in fig.˜9 (right) provides a more granular view, showing balanced accuracy across LoRA ranks on clean vs. spurious test sets.
@@ A.7 Aggresive SSTI: Robustness Through LoRA Rank
Meta-LLama-3.2-3B 1.06 0.68 0.66 0.64
A.8 Trends Across Locations
Table 11: The full table from table˜3. Accuracy degradation (↓, in percentage points) across two perturbation dimensions—injection location and token type—for snowflake-arctic-embed-l on the IMDB dataset. Results are shown for both Light and Aggressive SSTI (with 50% samples injected). An outlier for the light SSTI trend with date tokens, but is consistent across locations. Becomes consistent with the light SSTI trend: higher rank amplifies susceptibility for other token types, for date and HTML tokens. Fully consistent for aggressive SSTI: high rank improves robustness. For all cases, SSTI controls the behavior of the model.
SSTI Rank Injection Location Token Type
Beg. End Rand Date Country HTML
Light 1 4.14 4.21 4.24 4.21 0.67 0.74
16 4.14 4.07 4.09 4.07 2.07 1.79
32 4.02 3.82 3.91 3.82 2.91 2.45
64 3.80 3.62 3.59 3.62 3.00 2.84
Agg. 1 11.64 11.54 11.66 11.54 8.25 9.91
16 4.62 4.58 4.58 4.58 4.40 4.72
32 4.35 4.25 4.36 4.25 4.16 4.54
64 4.09 3.95 4.03 3.95 3.92 4.26
@@ A.8 Trends Across Locations
(b) Difference in balanced accuracy (↓) between spurious and clean test sets across LoRA ranks, under aggressive SSTI. Each curve corresponds to a different LoRA rank. LoRA rank amplifies resistance to spurious correlations when injection occurs at the end of samples.
(c) Difference in balanced accuracy (↓) between spurious and clean test sets across LoRA ranks, under aggressive SSTI. Each curve corresponds to a different LoRA rank. LoRA rank amplifies resistance to spurious correlations when injection occurs at at the beginning of the samples.
@@
A.9 Other Spurious Token Types
@@
A.10 Impact of token diversity
In this section, we look at how token diversity impacts our results. We ablate on date tokens due to the higher ceiling of unique tokens available, allowing us to test a wider range of token diversity that would be possible with html or countries.
@@ A.10 Impact of token diversity
A.11 Full finetuning
In this section, we conducted some full finetuning (without LoRA) experiments, to see if SSTI, also impacts an LLM finetuned through regular finetuning. We found that SSTI still has an impact on accuracy degradation during full finetuning of a pretrained model (as seen below table˜12).
Table 12: Difference in balanced accuracy between spurious and clean evaluation sets (accuracy degradation in $p$) for regular finetuning on IMDB. SSTI controls model behavior during regular finetuning also.
Dataset Model Accuracy Degradation (pp)
Full finetuning
IMDB Snowflake-arctic-embed-xs 4.61
Snowflake-arctic-embed-l 4.31
OpenELM-270M 1.46
OpenELM-3B 14.79
Meta-LLama-3.2-3B 6.23
A.12 Using a LoRA Variation
In this section, we look to see if the impact from SSTI are maintained when training with a variation of LoRA. We decided to run a small ablation by finetuning our two smallest models Snowflake-arctic-embed-xs and Snowflake-arctic-embed-l with DoRA (Liu et al., 2024a) under the same training conditions and parameters as the LoRA experiments.
Table 13: Snowflake-arctic-embed-xs and Snowflake-arctic-embed-l with aggressive SSTI, finetuned using DoRA. SSTI continues to generally impact models, and follows general trend seen with LoRA: in most cases, accuracy degradation (↓) with rank.
Model Dataset Accuracy Degradation (pp)
1 16 32 64
Snowflake-arctic-embed-xs IMDB 14.34 6.66 6.16 6.18
Financial Classification 0.12 4.98 4.98 4.48
Bias in Bios 0.00 0.42 0.42 0.54
Common Sense 6.82 10.04 10.04 10.04
Snowflake-arctic-embed-l IMDB 6.32 4.46 4.03 3.81
Financial Classification 5.10 5.10 4.48 4.48
Bias in Bios 0.44 0.84 0.89 0.90
Common Sense 10.04 9.96 9.73 9.33
Table 14: Snowflake-arctic-embed-xs and Snowflake-arctic-embed-l with light SSTI, finetuned using DoRA. SSTI continues to generally manipulate models.
Model Dataset Accuracy Degradation (pp)
1 16 32 64
Snowflake-arctic-embed-xs IMDB 2.23 6.50 6.26 5.74
Financial Classification 0.00 4.85 4.48 4.48
Bias in Bios 0.00 0.03 0.03 0.03
Common Sense 4.31 10.04 10.04 10.04
Snowflake-arctic-embed-l IMDB 5.34 3.94 3.80 3.60
Financial Classification 4.98 4.48 3.98 4.42
Bias in Bios 0.02 0.10 0.13 0.21
Common Sense 10.04 9.96 9.41 9.29
A.13 Further Examples for Recognizing SSTI